# Privacy-Preserving Distributed Link Predictions Among Peers in Online Classrooms Using Federated Learning

### v1

@@ 3.1 Social Learning Network (SLN) Structure
Figure 1: SLNs illustrating students’ connections formed at two subsequent time instants, where, $S_{1}$, $S_{2}$, etc., represent individual students. (a) Out of 10 students, four stay unconnected by the end of $t-1$, and (b) besides existing connections, new links were formed by the end of $t$ that can be attributed to course structure and student interactions.
@@ 3.1.2 Temporal Representation of SLNs
The MOOC datasets (i.e., ‘algo’, ‘comp’, ‘shake’, and ‘ml’) are static longitudinal data logs capturing interactions between online student pairs, with links formed cumulatively over a certain period. However, since explicit timeline information was not provided, we imposed a temporal structure by assuming the absence of links among certain pairs at an earlier time, thereby reflecting the potential for those links to form later. To emulate this temporal evolution, we created a modified version of the original datasets representing time step $t-1$ by randomly selecting 20% of the data (i.e., student pairs) and removing any existing links between the corresponding student pairs. The assumption is that the removed links would emerge at time step $t$, as reflected in the original datasets. Using this setup, we employed the data at time step $t-1$ to predict link formations at time step $t$. Nonetheless, the ‘csc’ dataset contains student interaction data recorded continuously throughout the semester. For this dataset, we captured the first instance of link formation between student pairs. Subsequently, we used the SLN topology at week 10 (i.e., time $t-1$) to predict the SLN topology — i.e., the links among students — in weeks 11, 12, 13, and 14 (i.e., time $t$). To evaluate our prediction performance, we partitioned the data into training and test sets, allocating 80% of the student pairs to the training set (i.e., 80% of student pairs at time $t-1$) and the remaining 20% to the test set.
@@ 3.1.3 Feature Description
$\mathcal{A}_{uv}=\sum_{S_{n}\in\Gamma_{u}\cap\Gamma_{v}}\frac{1}{\log|\Gamma_{n}|},\vspace{-.5mm}$ (2)
@@
$\mathcal{R}_{uv}=\sum_{S_{n}\in\Gamma_{u}\cap\Gamma_{v}}\frac{1}{|\Gamma_{n}|}.\vspace{-.5mm}$ (3)
@@
$\mathcal{P}_{uv}=|\Gamma_{u}|\cdot|\Gamma_{v}|.\vspace{-.5mm}$ (4)
@@
$\mathcal{C}_{uv}=\frac{|\Gamma_{u}\cap\Gamma_{v}|}{\sqrt{|\Gamma_{u}|\cdot|\Gamma_{v}|}},\vspace{-.5mm}$ (5)
@@
$\mathcal{D}_{uv}=\frac{2\cdot|\Gamma_{u}\cap\Gamma_{v}|}{|\Gamma_{u}|+|\Gamma_{v}|},\vspace{-.5mm}$ (6)
@@ 3.3 Federated Learning (FL) Architecture
$\mathbf{w}_{i}^{(k),e}=\mathbf{w}_{i}^{(k),e-1}-\eta\,\widetilde{\nabla}\mathcal{L}_{i}\big{(}\mathbf{w}_{i}^{(k),e-1}\big{)},\vspace{-.2mm}$ (7)
where $e\in\{1,\cdots,e_{\text{max}}\}$ indicates the SGD iteration count, $\eta$ denotes the learning rate, and ~$\widetilde{\nabla}\mathcal{L}_{i}(\mathbf{w}_{i}^{(k),e-1})$ is the stochastic approximation of the gradient of the local loss function, computed as ~$\widetilde{\nabla}\mathcal{L}_{i}(\mathbf{w}_{i}^{(k),e-1})=\frac{1}{|\mathcal{B}_{i}^{(k),e}|}\sum_{d\in\mathcal{B}_{i}^{(k),e}}\nabla\ell(\mathbf{w}_{i}^{(k),e-1};d),$ where $\mathcal{B}_{i}^{(k),e}\subseteq\mathcal{D}_{i}$ denotes a randomly selected mini-batch of data from the local dataset, and $|\mathcal{B}_{i}^{(k),e}|$ is its size. After completing the local model training, each client sends its latest local model parameters $\mathbf{w}_{i}^{(k),e_{\text{max}}}$ to the central server. It is worth noting that while the structure of the local models is consistent with that of the global model (i.e., all models share the same CNN architecture as in centralized training), the key difference lies in their trained parameters, which reflect the variations in local datasets and the differing patterns of student interactions they represent.
@@
$\mathbf{w}^{(k+1)}=\sum_{i\in\mathcal{I}}\frac{\big{|}\mathcal{D}_{i}\big{|}\mathbf{w}_{i}^{(k),e_{\text{max}}}}{|\mathcal{D}|},\vspace{-.2mm}$ (8)
@@ 3.5 Fairness Measures
We consider two widely adopted fairness measures in the literature: equalized odds and equalized opportunity [30]. Both measures are built upon quantifiable metrics — namely the true positive rate (TPR), which assesses the model’s ability to correctly identify students who form links, and the false positive rate (FPR), which evaluates how often students not connected are incorrectly classified as link-forming. Let $d$ denote an arbitrary data point (i.e., a student pair), $\widehat{Y}(d)$ the predicted label, and ${Y}(d)$ the true label (i.e., 00 or $1$ depending on the existence of an edge) of the data point. In this study, our goal is to assess the fairness of the model across various classrooms/SLNs (e.g., ensuring the FL global model treats students from different classrooms equally well). In this context, the equalized odds criterion is satisfied when both of the following probabilistic conditions hold for every pair of SLNs/clients $i,i^{\prime}\in\mathcal{I}$. (Condition 1) Equality of TPRs: $\mathbb{P}\big{(}\widehat{Y}(d)=1\big{|}Y(d)=1,d\in\mathcal{D}_{i}\big{)}=\mathbb{P}\big{(}\widehat{Y}(d)=1\big{|}Y(d)=1,d\in\mathcal{D}_{i^{\prime}}\big{)}$, (Condition 2) Equality of FPRs: $\mathbb{P}\big{(}\widehat{Y}(d)=1\big{|}Y(d)=0,d\in\mathcal{D}_{i}\big{)}=\mathbb{P}\big{(}\widehat{Y}(d)=1\big{|}Y(d)=0,d\in\mathcal{D}_{i^{\prime}}\big{)}$. In essence, equalized odds is achieved when both the TPR (i.e., Condition 1) and FPR (i.e., Condition 2) are equal across clients/SLNs. Further, equal opportunity is a relaxed version of equalized odds, requiring only the equality of TPRs (i.e., the satisfaction of Condition 1) across clients. In our subsequent simulations, we evaluate the fairness of the models by examining how closely their TPRs and FPRs adhere to these fairness criteria, thereby assessing the extent to which the models satisfy equalized odds and equal opportunity.
@@ 3.6 Model Explainability
$\varphi_{{}_{f}}=\sum_{\mathcal{F}\subseteq\mathcal{N}\setminus\{f\}}\frac{\big{|}\mathcal{F}\big{|}!(n-\big{|}\mathcal{F}\big{|}-1)!}{n!}\Big{(}v(\mathcal{F}\cup\{f\})-v(\mathcal{F})\Big{)},$ (9)
@@ 4 Results & Discussions
comp Loss $0.2\pm 0.01$ $0.22\pm 0.2$ $0.18\pm 0.003$ $0.21\pm 0.001$ $\textbf{0.18}\pm\textbf{0.002}$
@@ 4.2 Fairness Evaluation
We first focus on the TPR values in Table 3 and observe a higher consistency across the TPR values for FedALA compared to both centralized and FedAvg methods. This implies that FedALA adheres to the equal opportunity measure more closely. To quantify this finding, we compute the equal opportunity difference (EOD) metric [25], which measures the difference between TPR values across the classrooms, referred to as the range of TPR. A smaller range of TPR is desirable as it reflects more consistent positive predictions across different student pairs belonging to different SLNs, contributing to model fairness — the most fair model would have the range of 00. We depict the range of TPR of the methods in Figure 3(a), which verifies that FedALA achieves the lower TPR range, implying its higher adherence to equal opportunity measure. On the other hand, inspecting the FPR values in Table 3 there seem to be no conclusive results: the range (the difference between the maximum and minimum value computed over the SLNs) of FPR for FedALA is on par with centralized but is worse than FedAvg as also observed in Figure 3(b). So, a concrete conclusion on the equal odds measure cannot be drawn from the results. In summary, by exhibiting a better TPR consistency performance in terms of EOD, FedALA outperforms the other methods in terms of model fairness. Also, the results open a door to future studies on improving the adherence of FL models to the equal odds measure in SLNs.

### v2

@@ 3.1 Social Learning Network (SLN) Structure
Figure 1: SLNs illustrating students’ connections formed at two subsequent time instants, where, $S_{1}$, $S_{2}$, etc., represent individual students. (a) Out of 10 students, four stay unconnected by the end of t1t-1, and (b) besides existing connections, new links were formed by the end of $t$ that can be attributed to course structure and student interactions.
@@ 3.1.2 Temporal Representation of SLNs
The MOOC datasets (i.e., ‘algo’, ‘comp’, ‘shake’, and ‘ml’) are static longitudinal data logs capturing interactions between online student pairs, with links formed cumulatively over a certain period. However, since explicit timeline information was not provided, we imposed a temporal structure by assuming the absence of links among certain pairs at an earlier time, thereby reflecting the potential for those links to form later. To emulate this temporal evolution, we created a modified version of the original datasets representing time step t1t-1 by randomly selecting 20% of the data (i.e., student pairs) and removing any existing links between the corresponding student pairs. The assumption is that the removed links would emerge at time step $t$, as reflected in the original datasets. Using this setup, we employed the data at time step t1t-1 to predict link formations at time step $t$. Nonetheless, the ‘csc’ dataset contains student interaction data recorded continuously throughout the semester. For this dataset, we captured the first instance of link formation between student pairs. Subsequently, we used the SLN topology at week 10 (i.e., time t1t-1) to predict the SLN topology — i.e., the links among students — in weeks 11, 12, 13, and 14 (i.e., time $t$). To evaluate our prediction performance, we partitioned the data into training and test sets, allocating 80% of the student pairs to the training set (i.e., 80% of student pairs at time t1t-1) and the remaining 20% to the test set.
@@ 3.1.3 Feature Description
$\mathcal{A}_{uv}=\sum_{S_{n}\in\Gamma_{u}\cap\Gamma_{v}}\frac{1}{\log|\Gamma_{n}|},$ (2)
@@
$\mathcal{R}_{uv}=\sum_{S_{n}\in\Gamma_{u}\cap\Gamma_{v}}\frac{1}{|\Gamma_{n}|}.$ (3)
@@
$\mathcal{P}_{uv}=|\Gamma_{u}|\cdot|\Gamma_{v}|.$ (4)
@@
$\mathcal{C}_{uv}=\frac{|\Gamma_{u}\cap\Gamma_{v}|}{\sqrt{|\Gamma_{u}|\cdot|\Gamma_{v}|}},$ (5)
@@
$\mathcal{D}_{uv}=\frac{2\cdot|\Gamma_{u}\cap\Gamma_{v}|}{|\Gamma_{u}|+|\Gamma_{v}|},$ (6)
@@ 3.3 Federated Learning (FL) Architecture
$\mathbf{w}_{i}^{(k),e}=\mathbf{w}_{i}^{(k),e-1}-\eta\,\widetilde{\nabla}\mathcal{L}_{i}\big(\mathbf{w}_{i}^{(k),e-1}\big),$ (7)
where $e\in\{1,\cdots,e_{\text{max}}\}$ indicates the SGD iteration count, $\eta$ denotes the learning rate, and $\widetilde{\nabla}\mathcal{L}_{i}(\mathbf{w}_{i}^{(k),e-1})$ is the stochastic approximation of the gradient of the local loss function, computed as $\widetilde{\nabla}\mathcal{L}_{i}(\mathbf{w}_{i}^{(k),e-1})=\frac{1}{|\mathcal{B}_{i}^{(k),e}|}\sum_{d\in\mathcal{B}_{i}^{(k),e}}\nabla\ell(\mathbf{w}_{i}^{(k),e-1};d),$ where $\mathcal{B}_{i}^{(k),e}\subseteq\mathcal{D}_{i}$ denotes a randomly selected mini-batch of data from the local dataset, and $|\mathcal{B}_{i}^{(k),e}|$ is its size. After completing the local model training, each client sends its latest local model parameters $\mathbf{w}_{i}^{(k),e_{\text{max}}}$ to the central server. It is worth noting that while the structure of the local models is consistent with that of the global model (i.e., all models share the same CNN architecture as in centralized training), the key difference lies in their trained parameters, which reflect the variations in local datasets and the differing patterns of student interactions they represent.
@@
$\mathbf{w}^{(k+1)}=\sum_{i\in\mathcal{I}}\frac{\big|\mathcal{D}_{i}\big|\mathbf{w}_{i}^{(k),e_{\text{max}}}}{|\mathcal{D}|},$ (8)
@@ 3.5 Fairness Measures
We consider two widely adopted fairness measures in the literature: equalized odds and equalized opportunity [30]. Both measures are built upon quantifiable metrics — namely the true positive rate (TPR), which assesses the model’s ability to correctly identify students who form links, and the false positive rate (FPR), which evaluates how often students not connected are incorrectly classified as link-forming. Let $d$ denote an arbitrary data point (i.e., a student pair), $\widehat{Y}(d)$ the predicted label, and Y(d){Y}(d) the true label (i.e., 0 or $1$ depending on the existence of an edge) of the data point. In this study, our goal is to assess the fairness of the model across various classrooms/SLNs (e.g., ensuring the FL global model treats students from different classrooms equally well). In this context, the equalized odds criterion is satisfied when both of the following probabilistic conditions hold for every pair of SLNs/clients $i,i^{\prime}\in\mathcal{I}$. (Condition 1) Equality of TPRs: $\mathbb{P}\big(\widehat{Y}(d)=1\big|Y(d)=1,d\in\mathcal{D}_{i}\big)=\mathbb{P}\big(\widehat{Y}(d)=1\big|Y(d)=1,d\in\mathcal{D}_{i^{\prime}}\big)$, (Condition 2) Equality of FPRs: $\mathbb{P}\big(\widehat{Y}(d)=1\big|Y(d)=0,d\in\mathcal{D}_{i}\big)=\mathbb{P}\big(\widehat{Y}(d)=1\big|Y(d)=0,d\in\mathcal{D}_{i^{\prime}}\big)$. In essence, equalized odds is achieved when both the TPR (i.e., Condition 1) and FPR (i.e., Condition 2) are equal across clients/SLNs. Further, equal opportunity is a relaxed version of equalized odds, requiring only the equality of TPRs (i.e., the satisfaction of Condition 1) across clients. In our subsequent simulations, we evaluate the fairness of the models by examining how closely their TPRs and FPRs adhere to these fairness criteria, thereby assessing the extent to which the models satisfy equalized odds and equal opportunity.
@@ 3.6 Model Explainability
$\hskip-17.07164pt\varphi_{{}_{f}}=\sum_{\mathcal{F}\subseteq\mathcal{N}\setminus\{f\}}\hskip-5.69054pt\frac{\big|\mathcal{F}\big|!(n-\big|\mathcal{F}\big|-1)!}{n!}\Big(v(\mathcal{F}\cup\{f\})-v(\mathcal{F})\Big),\hskip-8.53581pt$ (9)
@@ 4 Results & Discussions
comp Loss $\pm 0.01 0.22±0$$2\pm 0.2$ $0.18\pm 0.003$ $0.21\pm 0.001$ $\textbf{0.18}\pm\textbf{0.002}$
@@ 4.2 Fairness Evaluation
The range (difference between the highest and lowest) of TPR values across all classrooms/SLNs.
@@
The range (difference between the highest and lowest) of FPR values across all classrooms/SLNs.
@@
We first focus on the TPR values in Table 3 and observe a higher consistency across the TPR values for FedALA compared to both centralized and FedAvg methods. This implies that FedALA adheres to the equal opportunity measure more closely. To quantify this finding, we compute the equal opportunity difference (EOD) metric [25], which measures the difference between TPR values across the classrooms, referred to as the range of TPR. A smaller range of TPR is desirable as it reflects more consistent positive predictions across different student pairs belonging to different SLNs, contributing to model fairness — the most fair model would have the range of 0. We depict the range of TPR of the methods in Figure 3(a), which verifies that FedALA achieves the lower TPR range, implying its higher adherence to equal opportunity measure. On the other hand, inspecting the FPR values in Table 3 there seem to be no conclusive results: the range (the difference between the maximum and minimum value computed over the SLNs) of FPR for FedALA is on par with centralized but is worse than FedAvg as also observed in Figure 3(b). So, a concrete conclusion on the equal odds measure cannot be drawn from the results. In summary, by exhibiting a better TPR consistency performance in terms of EOD, FedALA outperforms the other methods in terms of model fairness. Also, the results open a door to future studies on improving the adherence of FL models to the equal odds measure in SLNs.